OpenStudy (anonymous):

What are the solutions to the equation x2 + 3 = 124? A.–3 and 3 B.–10 and 10 C.–11 and 11 D.The equation has no solutions.

OpenStudy (anonymous):

Please explain

OpenStudy (saifoo.khan):

You missed an x.

OpenStudy (saifoo.khan):

x2 + 3x = 124?

OpenStudy (anonymous):

No i didn't thats what the equation says

OpenStudy (saifoo.khan):

x2 + 3 = 124 x^2 +3 - 124 = 0 x^2 - 121 = 0 NOT it's differences of squares. use the formula, (a^2 - b^2 ) = (a+b)(a-b) x^2 - 121 =0 (x-11)(x+11) = 0

OpenStudy (anonymous):

so it doesnt have an answer??

OpenStudy (anonymous):

x2 + 3 = 124 Then x*2 = 121 and x= 60,5 Or did you mean x^2? Then it's x^2-121= 0 = (x-11) (x+11) = 0 X=-11 and +11